esai5648:

The Hu family goes out for lunch, and the price of the meal is $53. The sales tax on the meal is 7%, and the family also leaves a 20% tip on the pre-tax amount. What is the total cost of the meal?

jacob28:

if ders choices dat would help

jacob28:

but wat...

Hoodmemes:

We gon find out these tax sales first, so the tax sells is 7% so we wanna multiply the cost of the meal by that after turning the tax sells into a decimal \[7 \div100=0.07\] now multiply \[0.07\times 53=3.71\] So,your total tax sells is 3.71 Now we find the total tip amount by multiply price of the meal by tip rate but we need to make the tip a decimal\[20\div 100=0.20\] Now multiply \[0.20\times 53=10.60\] Now to find the total cost u add \[53+10.60+3.71=(blank)\]

Hoodmemes:

(do u understand?)

Hoodmemes:

np

Hoodmemes:

So what do u think your final answer is?